An automobile tire has a volume of 1000 cm³. It is inflated with carbon dioxide to pressure of 40 psig at 23°C.How many grams of carbon dioxide are contained in the tire?


1
Expert's answer
2022-01-11T09:49:02-0500

pV = nRT

p = 40 psig = 2.7218 atm

V = 1000 cm3 = 1 L

T = 23°C = (23 + 273 K) = 296 K

n = pV/RT = (2.7218 atm) x (1 L) / (0.08206 L atm K-1 mol-1) x (296 K) = 0.112 mol

n = m/M

m = 0.112 mol x 44 g/mol = 4.93 g